# Simulation as Supervision: Mechanistic Pretraining for Scientific Discovery

### v3

@@ 4.3 Realistic, Diverse Mechanistic Grounding is Essential for Generalization
4.4 SGNNs Generalize Across Domains and Tasks
On higher-dimensional forecasting, SGNNs maintained robust forecasting skill as the number of butterfly species increased from 4 to 16 (Figure 3A, right), outperforming task-specific neural networks, which failed entirely in high dimensions. By 16 species, SGNN retained positive forecasting skill above 19.0, while baseline neural nets fell below zero. SGNNs can enable high-dimensional forecasting where standard approaches break down.
4.4.2 SGNNs Boost Chemical Yield Prediction with Minimal Pretraining
@@ 4.4.3 SGNNs Accurately Infer the Source of Diffusion Cascades
4.5 SGNNs Enable Accurate Inference of Unobservable Parameters
@@
The SGNN, trained solely on synthetic data, produced $R_{0}$ estimates from early COVID-19 data that were more consistent with later estimates once reporting had improved. For example, early estimates for New York City placed $R_{0}$ from 1.0-2.5 [33], yet subsequent studies using more complete data revised the estimate upward to over 5 [34, 35]. The SGNN, using only data from the early February-March period, correctly produced a high $R_{0}$ estimate for NYC (6.14), consistent with later, more complete analyses, versus 2.35 for exponential growth fitting. A full table of $R_{0}$ estimates is provided in Appendix E.
These results demonstrate that simulation-grounded training enables supervised learning for scientific quantities that are unobservable in real data and highly sensitive to noise and modeling assumptions. SGNN’s robustness to data imperfections and its ability to implicitly infer complex transmission dynamics from short, noisy time series make it well-suited for real-time epidemic inference.
Figure 4: SGNNs accurately infer unobservable parameters and provide mechanistic interpretability via back-to-simulation attribution. (A) SGNNs infer a high reproduction number ($R_{0}=6.14$) for New York City from early COVID-19 case data (Feb–Mar 2020), aligning with estimates from more complete datasets. Traditional methods underestimated early transmission due to underreporting and simplifying assumptions. (B) SGNNs achieve significantly lower mean squared error (left) and mean percentage error (right) in $R_{0}$ estimation compared to maximum likelihood estimation (MLE) and exponential growth-based methods. (C) Back-to-simulation attribution retrieves the 50 most similar synthetic outbreaks to a real-world input and visualizes the distribution of their underlying mechanistic parameters. For Michigan COVID-19 mortality, the retrieved simulations align closely with true estimates for hospitalization rate, population size, asymptomatic transmission rate, and hospitalization fatality rate (dashed lines), confirming that the SGNN’s internal representations encode meaningful mechanistic structure.
@@ 4.6 Back-to-Simulation Attribution Enables Mechanistic Interpretability
Beyond predictive performance, SGNNs enable back-to-simulation attribution, allowing real-world inputs to be mechanistically interpreted by identifying which simulated regimes the model deems most similar. Let $f_{\phi}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ denote the trained SGNN encoder that maps input time series to a $d$-dimensional embedding space. For a real-world observation $x$, we retrieve simulations $\{x_{i}\}$ from the training set $\mathcal{D}_{\text{syn}}$ by computing cosine similarity between $f_{\phi}(x)$ and $f_{\phi}(x_{i})$, selecting the top-$k$ most similar embeddings.
@@
Figure 4 shows violin plots of four key parameters across the top-50 retrieved simulations: case-hospitalization rate, hospitalization fatality rate, asymptomatic probability, and population size. These retrieved values align closely with known or plausible values for COVID-19 in Michigan, capturing the approximate population size (median population 10.1m, versus 10.14 actual), hospitalization and death probabilities consistent with empirical estimates, and realistic asymptomatic proportions. This alignment confirms that the SGNN grounds its predictions in simulations that resemble real-world system dynamics.
@@ 5 Discussion
SGNNs provide a unified framework for incorporating mechanistic knowledge into machine learning through training data. By generating synthetic datasets from mechanistic simulators, this approach enables neural networks to learn both system dynamics and the observational processes that shape real-world data, including noise, underreporting, and delays.
@@
Central to this framework is the shift from functional constraints to structural priors. While existing hybrid modeling approaches (e.g., PINNs or UDEs) enforce domain knowledge by explicitly constraining a model’s derivative or loss function, simulation-grounded learning embeds scientific structure through examples. This allows the neural network to internalize the global manifold of physically plausible dynamics without being tied to a single, potentially incorrect functional form. Because simulations span diverse model structures and parameter regimes, SGNNs develop a form of ensemble robustness, generalizing effectively even when real-world dynamics deviate from any single training simulator.
Training on simulated data also provides ground-truth labels for quantities that are unobservable in real settings, like transmission rates or carrying capacities that lack direct measurements. This enables approaches like simulation-based inference (SBI), which use supervised learning for parameter inference. Our results show that various approaches are possible within this framework: from full Bayesian posterior estimation, established with SBI, to simpler methods like quantile regression, which can achieve accurate parameter inference with reduced computational complexity.
@@
Back-to-simulation attribution further enhances the utility of this framework by linking real-world inputs to training simulations the model considers similar. Unlike post-hoc feature attribution methods, this reveals which mechanistic regimes the model associates with observed patterns. For example, if a model retrieves simulations with high asymptomatic transmission when shown an outbreak with few reported cases, it suggests the model interprets the pattern as indicating substantial hidden spread. This provides actionable, process-level insight, allowing practitioners to validate mechanistic hypotheses against domain knowledge or use them to guide intervention strategies.
@@
Our results suggest that simulation-grounded learning can capture generalizable representations of scientific structure that extend beyond the training distribution. The success of this paradigm has been further validated through the development of Mantis [36], a simulation-grounded foundation model for infectious disease forecasting. By scaling the SGNN framework to a massive synthetic corpus, Mantis achieves performance competitive with state-of-the-art models across diseases and geographies entirely out-of-the-box, without requiring real-world fine-tuning. This underscores the potential for simulation-pretrained models to serve as inference engines for science, though substantial work remains to understand the theoretical bounds of such generalization.
@@
Despite these strengths, SGNNs have important limitations. First, performance depends fundamentally on simulation quality: if key mechanisms are absent from the training simulators, the model cannot learn them. While training on diverse simulator ensembles provides robustness to individual model misspecification, systematic gaps across all training simulations will propagate to predictions. Second, generating comprehensive simulation corpora can be computationally expensive, particularly for high-fidelity mechanistic models. Though this is a one-time cost and inference remains fast, it may be prohibitive for extremely expensive simulators. Third, the sim-to-real gap remains a challenge. Even with realistic noise models and observational artifacts, systematic differences between synthetic and real data can hurt performance. Also, we lack theoretical guarantees about when SGNNs will generalize. While our results show promising zero-shot transfer (e.g., human-to-human models predicting vector-borne disease), we cannot predict a priori which cross-domain generalizations will succeed. Finally, back-to-simulation attribution reveals what the model believes is happening mechanistically, but that can be wrong. Retrieved simulations reflect learned associations that, while often informative, can be incorrect. Understanding when simulation-grounded learning succeeds and when it fails remains an important direction for future work.
SGNNs demonstrate that mechanistic simulations can serve as effective training data across diverse scientific domains. By positioning simulation-grounded learning as a robust alternative to constraint-based hybrid methods, we establish a design principle for combining scientific theory with deep learning flexibility: use mathematical simulators to define a structural prior over the system dynamics, and let the data-driven model learn to perform inference within those bounds.
@@ A.3 Synthetic Data Generation for Suzuki-Miyaura Cross-Coupling
  • Expert-driven heuristics, such as:
@@ B.1 Model Architecture
To demonstrate the flexibility of the SGNN framework across tasks and domains, we implemented a range of neural architectures tailored to each problem type. We found that straightforward, task-appropriate models performed well out-of-the-box when trained with simulation-grounded supervision. This suggests that the SGNN training paradigm provides a strong inductive bias that reduces reliance on heavy architecture tuning. Accordingly, we selected architectures that align naturally with each task structure—e.g., sequence models for time series and graph models for networks—to ensure strong baseline performance without domain-specific feature engineering.
Time Series Forecasting. For epidemiological and ecological forecasting, we used two classes of models:
@@
Network Diffusion Source Identification. For source inference on partially observed diffusion cascades, we used a 4-layer Graph Attention Network (GAT) with residual connections, batch normalization, ReLU activations, and dropout. GATs are well-suited for this task because they learn to assign attention weights over neighboring nodes, enabling the model to focus on the most informative connections in a noisy or incomplete cascade.
@@
Each node’s input consists of its observed infection time (or a sentinel value if unobserved) concatenated with its Laplacian positional encoding (LapPE), derived from the first 16 nontrivial eigenvectors of the graph Laplacian. These encodings provide global structural context and allow the model to distinguish between nodes with identical local neighborhoods but different roles in the global network topology—critical for inferring cascade origins. The GAT was trained to output a per-node probability distribution over possible sources, using cross-entropy loss.
Chemical Yield Prediction. For regression on simulated chemical reaction yields, we used a hybrid wide-and-deep SGNN architecture tailored to discrete combinatorial inputs. The model is designed to capture both linear interactions and higher-order nonlinear effects among reaction components.
@@
  • Inputs: Each reaction is represented by five categorical variables: aryl halide, boronate, ligand, base, and solvent. Each component is embedded in a 128-dimensional learnable vector space, and the resulting embeddings are concatenated into a single 640-dimensional feature vector.
  • Wide pathway: A linear layer maps the concatenated embeddings directly to a scalar prediction. This “wide” component captures additive main effects and is designed to quickly learn interpretable marginal trends across the combinatorial design space.
  • Deep pathway: A parallel MLP first passes the concatenated embeddings through a non-linear interaction layer, followed by a series of hidden layers (768 → 768 → 384 → 128 → 64) with ReLU activations, batch normalization, and dropout. This deep component captures complex, nonlinear interactions between reaction components—such as synergy or incompatibility between specific reagents.
@@
  • Multi-head output: The deep representation is fed into three heads:
@@
      A main regression head (yield prediction), which is summed with the wide pathway output.
@@
      An auxiliary head, providing an independent yield prediction to encourage representation diversity.
@@
      A confidence head, which outputs a [0,1] scalar using a sigmoid activation to represent model uncertainty.
  • Ensembling: The final prediction is a learnable weighted ensemble of the main and auxiliary heads. Ensemble weights are learned during training via a softmax-normalized two-element parameter vector.
This architecture is trained entirely on synthetic reaction data generated from simulation-based rulesets. All models use mean squared error loss on yield values and are evaluated in a zero-shot fashion on real reactions not seen during pretraining. The SGNN framework allows the model to learn a latent representation of reactivity space grounded in mechanistic constraints, without requiring experimental supervision.

### v4

@@ 4.3 Realistic, Diverse Mechanistic Grounding is Essential for Generalization
4.4 SGNNs Boost Chemical Yield Regression with Minimal Pretraining
@@ 4.5.1 SGNNs Accurately Infer the Source of Diffusion Cascades
4.5.2 Estimating the Transmissibility of COVID-19
@@
The SGNN, trained solely on synthetic data, produced $R_{0}$ estimates from early COVID-19 data that were more consistent with later estimates once reporting had improved. For example, early estimates for New York City placed $R_{0}$ from 1.0-2.5 [12], yet subsequent studies using more complete data revised the estimate upward to over 5 [44, 19]. As shown in figure 4, the SGNN, using only data from the early February-March period, correctly produced a high $R_{0}$ estimate for NYC (6.14), consistent with later, more complete analyses, versus 2.35 for exponential growth fitting. A full table of $R_{0}$ estimates is provided in Appendix E.
These results demonstrate that simulation-grounded training enables supervised learning for scientific quantities that are unobservable in real data and highly sensitive to noise and modeling assumptions. While only one demonstration, the SGNN’s robustness to data imperfections and its ability to implicitly infer complex transmission dynamics from short, noisy time series suggest it is well-suited for real-time epidemic inference.
Figure 4: SGNNs accurately infer unobservable parameters and provide mechanistic interpretability via back-to-simulation attribution. (A) SGNNs infer a high reproduction number ($R_{0}=6.14$) for New York City from early COVID-19 case data (Feb–Mar 2020, shown in the plot), aligning with estimates from more complete datasets [19]. (B) SGNNs achieve substantially lower mean squared error (left) and mean percentage error (right) in $R_{0}$ estimation compared to maximum likelihood estimation (MLE) and exponential growth-based methods. Ground truth is known in this case since we are testing on simulated data. (C) Back-to-simulation attribution retrieves the 50 most similar synthetic outbreaks to a real-world input and visualizes the distribution of their underlying mechanistic parameters. For Michigan COVID-19 mortality, the retrieved simulations align closely with true estimates for hospitalization rate, population size, asymptomatic transmission rate, and hospitalization fatality rate (dashed lines) [17], confirming that the SGNN’s internal representations encode meaningful mechanistic structure.
@@ 4.6 Back-to-Simulation Attribution Enables Mechanistic Interpretability
Because SGNNs are based in mechanistic models, they enable a new approach to mechanistic interpretability, back-to-simulation attribution, wherein we identify which simulated regimes the model deems most similar to a given real-world input. Let $f_{\phi}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ denote the trained SGNN encoder that maps input time series to a $d$-dimensional embedding space. For a real-world observation $x$, we retrieve simulations $\{x_{i}\}$ from the training set $\mathcal{D}_{\text{syn}}$ by computing cosine similarity between $f_{\phi}(x)$ and $f_{\phi}(x_{i})$, selecting the top-$k$ most similar embeddings.
@@
Figure 4 shows violin plots of four key parameters across the top-50 retrieved simulations: case-hospitalization rate, hospitalization fatality rate, asymptomatic probability, and population size. These retrieved values align closely with known or plausible values for COVID-19 in Michigan, capturing the approximate population size (median population 10.1m, versus 10.14 actual), hospitalization and death probabilities consistent with empirical estimates, and realistic asymptomatic proportions [17]. This alignment confirms that the SGNN grounds its predictions in simulations that resemble real-world system dynamics.
@@ 5 Discussion
By generating synthetic datasets from mechanistic simulators, the SGNN framework enables neural networks to learn both system dynamics and the observational processes that shape real-world data, including noise, underreporting, and delays. Central to this framework is the shift from functional constraints to structural priors. While existing hybrid modeling approaches (e.g., PINNs or UDEs [34, 33]) enforce domain knowledge by explicitly constraining a model’s derivative or loss function, simulation-grounded learning embeds scientific structure through examples. This allows the neural network to internalize the global manifold of physically plausible dynamics without being tied to a single, potentially incorrect mechanistic model structure. Because simulations span diverse model structures and parameter regimes, SGNNs develop a form of ensemble robustness, generalizing effectively even when real-world dynamics deviate from any single training simulator.
Training on simulated data also provides ground-truth labels for quantities that are unobservable in real settings, such as transmission rates or carrying capacities that lack direct measurements. This enables the integration of Simulation-Based Inference (SBI) [14, 22, 9], which traditionally uses supervised learning for parameter inference, directly into the SGNN framework. While SBI often focuses on full Bayesian posterior estimation, our results demonstrate that SGNNs support a range of inference strategies. For example, our $R_{0}$ estimation experiments used direct quantile regression on synthetic outbreak trajectories with known ground-truth reproduction numbers, bypassing the need for explicit likelihood evaluation or posterior sampling while still producing calibrated uncertainty intervals. Similarly, the back-to-simulation attribution approach implicitly recovers distributional information over mechanistic parameters by retrieving the nearest synthetic neighbors in embedding space, providing a nonparametric alternative to formal posterior inference. Depending on the scientific need, one can move from full posterior estimation to these computationally efficient alternatives, all while maintaining the scientific grounding provided by the synthetic training corpus. In this sense, SGNNs serve as a general vehicle for SBI, allowing these techniques to be applied to complex forecasting and regression tasks where they were previously isolated.
@@
Beyond parameter inference, back-to-simulation attribution also provides a distinct form of mechanistic interpretability. Unlike post-hoc feature attribution methods that identify which input variables influenced a prediction, the retrieved simulations reveal which mechanistic regimes the model associates with observed data. For example, if a model retrieves simulations characterized by high asymptomatic transmission when shown an outbreak with few reported cases, it suggests the model has internally identified a regime of substantial hidden spread. This provides actionable, process-level insight, allowing practitioners to validate the model’s reasoning against domain knowledge and use those mechanistic insights to guide intervention strategies.
@@
Our results suggest that simulation-grounded learning can capture generalizable representations of scientific structure that extend beyond the training distribution. The success of this paradigm has since been further validated through the development of Mantis [11], a more advanced simulation-grounded foundation model for infectious disease forecasting. By scaling the SGNN framework to a massive synthetic corpus, Mantis achieves performance competitive with state-of-the-art models across diseases and geographies entirely out-of-the-box, without requiring real-world fine-tuning.
@@
Despite these strengths, SGNNs have important limitations that we plan to explore further in future work. First, performance depends fundamentally on simulation quality: if key mechanisms are absent from the training simulators, performance can be substantially degraded. While training on diverse simulator ensembles provides robustness to individual model misspecification, systematic gaps across all training simulations will propagate to predictions, underscoring the importance of careful choices around the mechanistic models and parameter ranges used for training. Even with realistic noise models and observational artifacts, we anticipate that systematic differences between synthetic and real data will hurt performance. Also, we lack theoretical guarantees about when SGNNs will generalize. While our results show promising zero-shot transfer (e.g., human-to-human models predicting vector-borne disease), we cannot predict a priori which cross-domain generalizations will succeed. Second, generating comprehensive simulation corpora can be computationally expensive, particularly for high-fidelity mechanistic models. Though this is a one-time cost and inference remains fast, it may be prohibitive for extremely expensive simulators. Finally, back-to-simulation attribution reveals what the model believes is happening mechanistically, but that can be wrong. Retrieved simulations reflect learned associations that, while often informative, can be incorrect and depend heavily on the simulation prior. Understanding when simulation-grounded learning succeeds and when it fails remains an important direction for future work.
6 Author Contributions
CD led the study and developed the code and initial draft of the manuscript. CH assisted with initial code development and RM developed figure visualizations. MCE, CD, CH, and RM all contributed substantial edits and revisions for the final manuscript.
@@ A.3 Synthetic Data Generation for Suzuki-Miyaura Cross-Coupling
  • Heuristics such as:
@@ B.1 Model Architecture
To demonstrate the flexibility of the SGNN framework across tasks and domains, we implemented a range of neural architectures tailored to each problem type. We found that straightforward, task-appropriate models performed well out-of-the-box when trained with simulation-grounded supervision. This suggests that the SGNN training framework reduces reliance on choosing a specific architecture. We selected architectures that align naturally with each task structure: e.g., sequence models for time series and graph models for networks.
Time Series Forecasting.
For epidemiological and ecological forecasting, we used two classes of models:
@@
Network Diffusion Source Identification.
To infer the origin of a diffusion cascade from noisy, partially observed data, we implemented a Graph Attention Network (GAT) [43] enhanced with spectral positional encodings. This architecture allows the model to leverage both the temporal dynamics of the cascade and the topology of the underlying network.
  • Input Representation and Laplacian PE: Each node $v$ is initialized with a feature vector $\mathbf{h}_{v}=[t_{v}||\mathbf{p}_{v}]$, where $t_{v}$ is the observed infection time (or a sentinel value for unobserved nodes) and $\mathbf{p}_{v}\in\mathbb{R}^{16}$ is a Laplacian Positional Encoding (LapPE) [24]. To compute $\mathbf{p}_{v}$, we define the symmetric normalized Laplacian $\mathcal{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$. We perform an eigendecomposition of $\mathcal{L}$ and retain the first $k=16$ non-trivial eigenvectors (excluding the smallest eigenvalue $\lambda_{0}=0$). These spectral encodings act as low-frequency structural coordinates, enabling the GAT to distinguish between nodes with identical local neighborhoods but distinct global roles, such as bridge nodes versus peripheral nodes.
  • Graph Attention Backbone: The model consists of 4 GAT layers with multi-head attention ($M=4$ heads). For a node $i$, the output of a GAT layer is:
    $\mathbf{x}^{\prime}_{i}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{x}_{j}\right)$ (6)
    where $\alpha_{ij}$ are the learnable attention coefficients computed via a shared self-attention mechanism on the node features. Each layer utilizes a hidden dimensionality of 256 ($64\times 4$ heads), followed by Batch Normalization, ReLU activation, and Dropout ($p=0.3$).
  • Residual Connectivity: To facilitate the training of a deep graph architecture and preserve signal from the temporal inputs, we implement additive skip connections. For each layer $l$, the update rule is $\mathbf{x}^{(l)}=\text{ReLU}(\text{GAT}(\mathbf{x}^{(l-1)}))+\mathbf{x}^{(l-1)}$, ensuring that the mechanistic signal from the infection times is propagated through the network.
  • Output Head and Inference: The final node embeddings are passed through a linear layer to produce a scalar logit for each node. These are reshaped into a probability distribution over the $N=1000$ nodes in the graph using a Softmax activation. The model is trained using Cross-Entropy loss to maximize the likelihood assigned to the true source node.
@@
Each node’s input consists of its observed infection time (or a sentinel value if unobserved) concatenated with its Laplacian positional encoding (LapPE) [24], derived from the first 16 nontrivial eigenvectors of the graph Laplacian. These encodings provide global structural context and allow the model to distinguish between nodes with identical local neighborhoods but different roles in the global network topology—critical for inferring cascade origins. The GAT was trained to output a per-node probability distribution over possible sources, using cross-entropy loss.
Chemical Yield Prediction.
For regression on simulated chemical reaction yields, we implemented a hybrid wide-and-deep SGNN architecture [7] specifically tailored to capture the discrete combinatorial nature of organic synthesis. This architecture simultaneously models global additive trends and local, high-order mechanistic interactions between reagents.
@@
  • Inputs and Embedding Layer: Each reaction is defined by five categorical variables: aryl halide, boronate, ligand, base, and solvent. Each component is mapped to a 128-dimensional learnable embedding. These are concatenated to form a 640-dimensional base feature vector $\mathbf{x}_{\text{emb}}$.
  • Wide Pathway: A linear transformation maps $\mathbf{x}_{\text{emb}}$ directly to a scalar. This wide component acts as a generalized linear model, capturing the additive marginal effects of each reagent and providing a stable baseline for the predicted yield.
  • Deep Pathway and Interaction Layer: The deep component begins with a 128-unit ”interaction” layer with ReLU activation that operates on the concatenated embeddings. The output is then concatenated with the original $\mathbf{x}_{\text{emb}}$ to form a 768-dimensional input for a multi-layer perceptron (MLP). The MLP consists of a series of dense layers (768 $\rightarrow$ 768 $\rightarrow$ 384 $\rightarrow$ 128 $\rightarrow$ 64) utilizing ReLU activations, Batch Normalization to stabilize the learning of synthetic distributions, and Dropout ($p\in[0.10,0.15]$) to prevent the model from memorizing specific simulation artifacts.
@@
  • Multi-Head Output and Ensembling: The final 64-dimensional latent representation is passed to three specialized heads:
@@
      Main Regression Head: Predicts a yield value that is summed with the wide pathway output to produce the primary prediction $\hat{y}_{m}$.
@@
      Auxiliary Head: A 2-layer MLP (64 $\rightarrow$ 32 $\rightarrow$ 1) that produces a secondary yield estimate $\hat{y}_{a}$, encouraging the deep feature extractor to maintain a diverse representation of the reactivity space.
@@
      Confidence Head: An MLP (64 $\rightarrow$ 32 $\rightarrow$ 1) with a sigmoid activation that outputs a scalar $\hat{c}\in[0,1]$. This head learns to weight the absolute residuals, effectively acting as a proxy for the model’s epistemic uncertainty regarding specific reagent combinations.
  • Dynamic Ensembling: The final model output $\hat{y}_{\text{ens}}$ is a convex combination of the main and auxiliary heads: $\hat{y}_{\text{ens}}=w_{m}\hat{y}_{m}+w_{a}\hat{y}_{a}$. The weights $\mathbf{w}$ are derived from a two-element parameter vector passed through a softmax function, allowing the model to learn the optimal balance between the heads during pretraining.
This architecture is initially trained on the large synthetic corpus to internalize the underlying chemical rules. By leveraging the wide-and-deep structure, the SGNN learns a latent manifold of chemical reactivity that generalizes from structured simulations to sparse, noisy experimental data.